# Black Hole Survival Guide: Searching for Stars in the Galactic Center That Endure Partial Tidal Disruption

### v1

@@ Abstract
Once per $\approx 10^{4}$$10^{5}$ years, an unlucky star may experience a close encounter with a supermassive black hole (SMBH), partially or fully tearing apart the star in an exceedingly brief, bright interaction called a tidal disruption event (TDE). Remnants of partial TDEs are expected to be plentiful in our Galactic Center, where at least six unexplained, diffuse, star-like “G objects” have already been detected which may have formed via interactions between stars and the SMBH. Using numerical simulations, this work aims to identify the characteristics of TDE remnants. We take 3D hydrodynamic FLASH models of partially disrupted stars and map them into the 1D stellar evolution code MESA to examine the properties of these remnants from tens to billions of years after the TDE. The remnants initially exhibit a brief, highly luminous phase, followed by an extended cooling period as they return to stable hydrogen burning. During the initial stage ($\lesssim 10^{4}$ yr) their luminosities increase by orders of magnitude, making them intriguing candidates to explain a fraction of the mysterious G objects. Notably, mild TDEs are the most common and result in the brightest remnants during this initial phase. However, most remnants exist in a long-lived stage where they are only modestly offset in temperature and luminosity compared to main-sequence stars of equivalent mass. Nonetheless, our results indicate remnants will sustain abnormal, metal-enriched envelopes that may be discernible through spectroscopic analysis. Identifying TDE survivors within the Milky Way could further illuminate some of the most gravitationally intense encounters in the Universe.
@@ 1 Introduction
Specifically, peculiar dust- and gas-enshrouded stellar objects, called G objects, have been observed in the Milky Way’s nuclear star cluster (S. Gillessen et al., 2012): six known G objects are presented in A. Ciurlo et al. (2020). Their origins remain unknown, although several formation channels have been proposed in the literature (e.g., A. Burkert et al., 2012; M. Schartmann et al., 2012; J. Guillochon et al., 2014; A.-M. Madigan et al., 2017; J. E. Owen & D. N. C. Lin, 2023). The proposed formation channels include, for example, stellar mergers (e.g., G. Witzel et al., 2014; S. Prodan et al., 2015; A. P. Stephan et al., 2016, 2019; S. Wu et al., 2020) and stellar collisions (e.g., S. C. Rose et al., 2023). Here we investigate the viability of a third formation avenue: the puffy stellar remnant formed when a star survives a tidal encounter with the central SMBH.
@@
At the center of the Milky Way, the gravitational landscape is dominated by Sgr A*, a SMBH with a mass roughly four million times that of our Sun (e.g., A. M. Ghez et al., 2008). Approximately 10 million stars swarm this SMBH in tightly packed orbits within a fraction of a parsec of the Galactic Center, an environment that is $\lesssim 10^{8}$ times denser than our Solar neighborhood (e.g., J. R. Lu et al., 2013). Each star within this region traces out an intricate orbit under the combined influence of the SMBH and all other stars. The tidal radius $R_{\tau}=R_{\ast}(M_{\rm bh}/M_{\ast})^{1/3}$ is defined as the boundary of closest approach, beyond which point the SMBH’s gravity overcomes the star’s self gravity and the SMBH begins to unbind the star and siphon stellar mass. Here $M_{\rm bh}$ is the mass of the SMBH and $M_{\ast}$ and $R_{\ast}$ are the stellar mass and radius, respectively. If a star wanders past this critical distance, a tidal disruption event (TDE) occurs, as the star is fiercely ripped apart by the SMBH’s tidal field (e.g., J. G. Hills, 1975; M. J. Rees, 1988; T. Alexander, 1999; J. Magorrian & S. Tremaine, 1999; J. Wang & D. Merritt, 2004).
To account for the distorted internal structure of the star during a TDE, a hydrodynamical description must be employed to effectively describe the remnant’s deformation and the exchange of energy and angular momentum during the encounter (C. R. Evans & C. S. Kochanek, 1989; P. B. Ivanov & I. D. Novikov, 2001; A. Khokhlov et al., 1993; P. Laguna et al., 1993; E. Ramirez-Ruiz & S. Rosswog, 2009; M. MacLeod et al., 2016; J. A. P. Law-Smith et al., 2020a). After the dust has settled, new stellar models may be constructed for the remnants, often with highly unusual chemical compositions and physical conditions. The timescales for the remnants to regain their thermal equilibrium are vastly longer than the minutes or days needed for dynamical equilibrium to be restored, which warrants pivoting to one-dimensional stellar evolution simulations in order to capture the further evolution (e.g., F. R. N. Schneider et al., 2020; S. Wu et al., 2020).
@@
TDEs are important to study because they can illuminate previously quiescent SMBHs, revealing key properties about those SMBHs in the process and ultimately contributing to population demographics (e.g., J. Guillochon & E. Ramirez-Ruiz, 2013; C. Yang et al., 2017; J. L. Dai et al., 2021; B. Mockler et al., 2022; J. M. Miller et al., 2023). Tidal encounters between SMBHs and stars are also phenomenal opportunities to deepen our understanding of environments with intense gravity (e.g., J. Guillochon et al., 2009). The catch is, TDEs are rare. The prevalence of TDEs is heavily dependent on cosmic epoch, merger history, and galaxy properties, but a conservative estimate suggests an occurrence rate of approximately $10^{-4}$ to $10^{-5}$ TDEs per galaxy per year (T. Alexander, 1999; J. Magorrian & S. Tremaine, 1999; J. Wang & D. Merritt, 2004; N. C. Stone & B. D. Metzger, 2016; H. Pfister et al., 2021). In other words, to guarantee seeing a TDE in progress would require staring at the nucleus of the Milky Way for 100,000 years, or alternatively, searching 100,000 galaxies every day for a year. Most have chosen the latter route, utilizing large surveys of galaxies in hopes of chancing upon the transient flare of a TDE in a galaxy far, far away. At the time of this paper’s publication, fewer than one hundred candidate TDEs have been observed using this strategy. [^1] https://github.com/astrocatalogs/tidaldisruptions
In this Letter, we aim to ascertain what the remnants of these tidal encounters look like in our own backyard – the Galactic Center of the Milky Way. A significant portion of TDEs do not result in total annihilation of the star, leaving behind a warped stellar remnant that survives and is ejected from the BH (N. C. Stone & B. D. Metzger, 2016). In the 13.6-billion-year history of the Milky Way, for a TDE occurrence rate of $\sim 10^{-5}\,\mathrm{yr}^{-1}$, there should have been time for approximately 100,000 such “survivor stars” to accumulate in our midst.
@@
Figure 1: Slices in 2D of the four TDE remnants studied in this paper, selected from the STARS Library (J. A. P. Law-Smith et al., 2020a). The left three columns show remnants that begin as a 1 $M_{\odot}$, 4.8 Gyr main sequence star before disruption, and the fourth displayed in the rightmost column begins as a 3 $M_{\odot}$, 0.3 Gyr main sequence star. Each is subjected to varying degrees of tidal disruption. The impact parameter ($\beta$) increases from left to right, indicating how deeply the star plunged past the SMBH’s tidal radius. Colorbars indicate temperature in Kelvin (top row) and diffusion time in years (bottom row). The left panels have width 10 $R_{\odot}$, while the rightmost column has width 22 $R_{\odot}$. Each panel is captured at $\approx 1000$ days after disruption.
@@
These remnant stars could bring insights into some of the most intense encounters with extreme gravity in the Universe. Here we investigate how TDE survivors evolve, from tens to billions of years after disruption, and we seek the fingerprints that will differentiate them from surrounding stars. Our research combines hydrodynamical and stellar evolutionary simulations to identify whether TDE survivors retain any signature characteristics that could clue us into their pasts. In our study, we capture the TDE survivor’s early ($>1000$ yr) evolution as well as the remaining $\gtrsim 100$ Myr of the remnant’s lifetime using open-source 3D hydrodynamical and 1D stellar evolution codes. Our results at late times are consistent with recent work investigating the long-term evolution ($>100$ Myr) of TDE survivors from M. Sharma et al. (2024). Our methods enable us to investigate the initial bright phase of the remnant and thereby quantify the remnant’s behavior during what may be the most opportune period for their detection as unusual stellar objects.
@@ 2 Numerical Framework
We aim to capture the details of the tidal disruption, which takes place over a few dynamical timescales of the star, as well as the remnant’s subsequent long-term evolution, lasting upwards of a few Gyr. To address this challenge, we have combined simulations of TDEs using the adaptive mesh hydrodynamic code FLASH (B. Fryxell et al., 2000) paired with stellar evolution models from the Modules for Experiments in Stellar Astrophysics code (MESA version 10398, B. Paxton et al. 2011, 2013, 2015, 2018) that explore the subsequent long-term evolution of the disrupted star. The hydrodynamic simulations are explained in Section 2.1, while the stellar evolution setup is detailed in Section 2.2.
@@ 2.1 Hydrodynamic FLASH models
In J. A. P. Law-Smith et al. (2020a), stellar models of varying initial masses and ages from MESA were mapped into 3D FLASH simulations (J. Guillochon et al., 2009; J. Guillochon & E. Ramirez-Ruiz, 2013), where they were tidally disrupted by a SMBH with mass $M_{\rm BH}=10^{6}\,M_{\odot}$. The results of these simulations are cataloged in the STARS Library (J. A. P. Law-Smith et al., 2020b). For our study of the long-term stellar evolution post-disruption, we select three disruptions at solar metallicity ($Z=0.0142$) of a star with initial mass $M_{i}=1\,M_{\odot}$ which were tidally disrupted at an age of 4.8 Gyr. Based on the known population of stars in the Galactic Center, we choose to primarily work with solar-mass stars to strike a middle ground between massive stars, which are rare, and low-mass stars, which are more abundant but are often fully disrupted, leaving no remnant. We selected middle-aged main sequence stars (MAMS) as our progenitors, as that is the evolutionary phase where stars spend the majority of their lives. We also model a fourth remnant from an initially $M_{i}=3\,M_{\odot}$ progenitor, disrupted at an age of 0.3 Gyr (MAMS).
@@
The models with a $M_{i}=1\,M_{\odot}$ progenitor underwent partial tidal disruption with varying values of $\beta=1$, 1.5, and 2. We define $\beta$ as the ratio of the tidal disruption radius ($R_{\tau}$) to the pericenter distance – a dimensionless parameter that characterizes the strength of the disruption. For instance, $\beta=1$ corresponds to the tidal radius, the point at which the SMBH’s tidal forces first begin to siphon mass from the star, and this was our lower threshold of “mild” partial disruption. For this progenitor, we select $\beta$ values distributed between 1 and 2 to best capture the possible range of disruption remnants. Above our upper threshold (“severe” disruption) of $\beta=2$ for the $M_{i}=1\,M_{\odot}$ progenitor, stars in the STARS Library are usually fully, or nearly fully, destroyed (see Figure 5 in J. A. P. Law-Smith et al., 2020a). This set of remnant models represents the range of outcomes for common stars that are likely to be disrupted, while also being likely to survive. The corresponding final masses of each TDE remnant model from $M_{i}=1\,M_{\odot}$ are $M_{f}=0.956\,~{}M_{\odot}$, $0.708\,~{}M_{\odot}$, and $0.386\,~{}M_{\odot}$ for $\beta=1$, 1.5, and 2 respectively. We refer to these models by their $\beta$ value and progenitor mass, e.g., the $\beta=1$, $M_{i}=1\,M_{\odot}$ remnant.
@@
Our fourth remnant is partially tidally disrupted with $\beta=4$ from a progenitor with $M_{i}=3\,M_{\odot}$. Post-disruption, the remnant mass is $M_{f}=1.06\,M_{\odot}$, nearly the same as the initial mass of our other models. Its final mass is also very similar to the post-disruption mass of the $\beta=1$, $M_{i}=1\,M_{\odot}$ remnant. As a result, the long-term evolution of the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant provides an interesting contrast to that of similar-mass models with different initial conditions.
@@
Figure 1 shows snapshots of the FLASH simulations of the four TDE remnants immediately post-disruption. The three panels on the left compare the three remnants from the same progenitor of $M_{i}=1\,M_{\odot}$. After the disruption, the least disrupted remnant ($\beta=1$) retains a dense, hot core with a high internal diffusion time, and a relatively short diffusion time in the envelope. This suggests the TDE primarily deposited energy close to the star’s surface without impacting much of the stellar interior. On the opposite end, the most disrupted remnant ($\beta=2$) is relatively cooler throughout with a diminished diffusion time at the core, indicating the TDE had a greater impact on the star’s internal structure. Energy deposited closer to the surface where the diffusion time is short will leak out first, powering the early-time luminosity of the remnant.
@@
On the right, we show snapshots from the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant, whose radial extent is a factor of $\sim 2$ larger than the remnants from a $M_{i}=1\,M_{\odot}$ progenitor. As this progenitor was disrupted at a very large $\beta$, the remnant also exhibits a short diffusion time throughout the core that signifies deep energy deposition from the TDE. Given the increased radial scale of these snapshots, we can see that the $\beta=4$, $M_{i}=3\,M_{\odot}$ remnant is much more diffuse, but still retains a relatively high core temperature due to its more massive progenitor.
@@
Figure 2: Luminosity, temperature and radius evolution of the three tidally disrupted remnants (red solid) with progenitor mass $M_{i}=1\,M_{\odot}$. These remnants are compared to mass-equivalent stars (blue dash-dotted) that evolve from the pre-main sequence (PMS) to the main sequence (MS) and up the red giant branch (RGB). The masses of the TDE remnants and their associated mass-equivalent stars are $0.956\,M_{\odot}$, $0.708\,M_{\odot}$, and $0.386\,M_{\odot}$ respectively from left to right. The start of core hydrogen fusion is denoted by star symbols for remnants and mass-equivalent stars. In addition, the evolution of an undisturbed $M=1\,M_{\odot}$ star (black dotted) is shown from the time of disruption onward in order to represent the comparative evolution of the progenitor had it not undergone the TDE. Immediately after disruption, the remnants appear larger, hotter, and more luminous than the progenitor, then shrink over the next several Myr. After reigniting core hydrogen fusion, they stabilize near the MS locations of their mass equivalent stars.

### v2

@@ Abstract
Once per $\approx 10^{4}$$10^{5}$ years, an unlucky star may experience a close encounter with a supermassive black hole (SMBH), partially or fully tearing apart the star in an exceedingly brief, bright interaction called a tidal disruption event (TDE). Remnants of partial TDEs are expected to be plentiful in our Galactic Center, where at least six unexplained, diffuse, star-like “G objects” have already been detected which may have formed via interactions between stars and the SMBH. Using numerical simulations, this work aims to identify the characteristics of TDE remnants. We take 3D hydrodynamic FLASH models of partially disrupted stars and map them into the 1D stellar evolution code MESA to examine the properties of these remnants from tens to billions of years after the TDE. The remnants initially exhibit a brief, highly luminous phase, followed by an extended cooling period as they return to stable hydrogen burning. During the initial stage ($\lesssim 10^{5}$ yr) their luminosities increase by orders of magnitude, making them intriguing candidates to explain a fraction of the mysterious G objects. Notably, mild TDEs are the most common and result in the brightest remnants during this initial phase. However, most remnants exist in a long-lived stage where they are only modestly offset in temperature and luminosity compared to main-sequence stars of equivalent mass. Nonetheless, our results indicate remnants will sustain abnormal, metal-enriched envelopes that may be discernible through spectroscopic analysis. Identifying TDE survivors within the Milky Way could further illuminate some of the most gravitationally intense encounters in the Universe.
@@ 1 Introduction
Specifically, peculiar dust- and gas-enshrouded stellar objects, called G objects, have been observed in the Milky Way’s nuclear star cluster (S. Gillessen et al., 2012; G. Witzel et al., 2014): six known G objects are presented in A. Ciurlo et al. (2020). Their origins remain unknown, although several formation channels have been proposed in the literature (e.g., A. Burkert et al., 2012; M. Schartmann et al., 2012; J. Guillochon et al., 2014; A.-M. Madigan et al., 2017; J. E. Owen & D. N. C. Lin, 2023). The proposed formation channels include, for example, stellar mergers (e.g., G. Witzel et al., 2014; S. Prodan et al., 2015; A. P. Stephan et al., 2016, 2019; S. Wu et al., 2020) and stellar collisions (e.g., S. C. Rose et al., 2023). Here we investigate the viability of a third formation avenue: the puffy stellar remnant formed when a star survives a tidal encounter with the central SMBH.
@@
At the center of the Milky Way, the gravitational landscape is dominated by Sgr A*, a SMBH with a mass roughly four million times that of our Sun (e.g., A. M. Ghez et al., 2008). Approximately 10 million stars swarm this SMBH in tightly packed orbits within a few parsecs of the Galactic Center, an environment that is $\lesssim 10^{8}$ times denser than our Solar neighborhood (e.g., J. R. Lu et al., 2013; R. Schödel et al., 2014, 2018). Each star within this region traces out an intricate orbit under the combined influence of the SMBH and all other stars. The tidal radius $R_{\tau}\equiv R_{\ast}(M_{\rm BH}/M_{\ast})^{1/3}$ is defined as the boundary of closest approach, beyond which point the SMBH’s gravity overcomes the star’s self gravity and the SMBH begins to unbind the star and siphon stellar mass. Here $M_{\rm BH}$ is the mass of the SMBH and $M_{\ast}$ and $R_{\ast}$ are the stellar mass and radius, respectively. If a star wanders past this critical distance, a tidal disruption event (TDE) occurs, as the star is fiercely ripped apart by the SMBH’s tidal field (e.g., J. G. Hills, 1975; M. J. Rees, 1988; T. Alexander, 1999; J. Magorrian & S. Tremaine, 1999; J. Wang & D. Merritt, 2004).
@@
TDEs are important to study because they can illuminate previously quiescent SMBHs, revealing key properties about those SMBHs in the process and ultimately contributing to population demographics (e.g., J. Guillochon & E. Ramirez-Ruiz, 2013; C. Yang et al., 2017; J. L. Dai et al., 2021; B. Mockler et al., 2022; J. M. Miller et al., 2023). Tidal encounters between SMBHs and stars are also phenomenal opportunities to deepen our understanding of environments with intense gravity (e.g., J. Guillochon et al., 2009). However, TDEs are rare events. The prevalence of TDEs is heavily dependent on cosmic epoch, merger history, and galaxy properties, but a conservative estimate suggests an occurrence rate of approximately $10^{-4}$ to $10^{-5}$ TDEs per galaxy per year (T. Alexander, 1999; J. Magorrian & S. Tremaine, 1999; J. Wang & D. Merritt, 2004; N. C. Stone & B. D. Metzger, 2016; H. Pfister et al., 2021). In other words, to guarantee seeing a TDE in progress would require staring at the nucleus of the Milky Way for 100,000 years, or alternatively, searching 100,000 galaxies every day for a year. Most have chosen the latter route, utilizing large surveys of galaxies in hopes of chancing upon the transient flare of a TDE in a galaxy far, far away. At the time of this paper’s publication, fewer than one hundred candidate TDEs have been observed using this strategy. [^1] https://github.com/astrocatalogs/tidaldisruptions
[…]
@@
Figure 1: Slices in 2D of the four TDE remnants studied in this paper, selected from the STARS Library (J. A. P. Law-Smith et al., 2020a). The left three columns show remnants that begin as a 1 $M_{\odot}$, 4.8 Gyr main sequence star before disruption, and the fourth displayed in the rightmost column begins as a 3 $M_{\odot}$, 0.3 Gyr main sequence star. Each is subjected to varying degrees of tidal disruption. The impact parameter ($\beta$) increases from left to right, indicating how deeply the star plunged past the SMBH’s tidal radius. Colorbars indicate temperature in Kelvin (top row) and diffusion time in years (bottom row). The left panels have width 10 $R_{\odot}$, while the rightmost column has width 22 $R_{\odot}$. All snapshots are captured at roughly 80 $t_{\rm dyn}$ after pericenter, corresponding to $\approx 38$ hr after pericenter for the left three columns and $\approx 124$ hr after pericenter for the right column.
To account for the distorted internal structure of the star during a TDE, a hydrodynamical description must be employed to effectively describe the remnant’s deformation and the exchange of energy and angular momentum during the encounter (C. R. Evans & C. S. Kochanek, 1989; P. B. Ivanov & I. D. Novikov, 2001; A. Khokhlov et al., 1993; P. Laguna et al., 1993; E. Ramirez-Ruiz & S. Rosswog, 2009; M. MacLeod et al., 2016; J. A. P. Law-Smith et al., 2020a). Then a new stellar model may be constructed for the remnant, often with a highly unusual chemical composition and physical conditions. The timescale for the remnant to regain its thermal equilibrium is vastly longer than the minutes or days needed for dynamical equilibrium to be restored, which warrants pivoting to one-dimensional stellar evolution simulations in order to capture the further evolution (e.g., F. R. N. Schneider et al., 2020; S. Wu et al., 2020).
@@
In our study, we utilize the open-source 3D hydrodynamical code FLASH (B. Fryxell et al., 2000) to simulate the TDE, then pivot to the 1D stellar evolution code MESA (B. Paxton et al., 2011, 2013, 2015, 2018) to capture the TDE survivor’s subsequent evolution. Our results at late times are consistent with recent work investigating the long-term evolution (>100>100 Myr) of TDE survivors from M. Sharma et al. (2024). What sets our study apart is that our methods enabled us to investigate the initial (>1000>1000 yr) bright phase of the remnant as well, and thereby quantify the remnant’s behavior during what may be the most opportune period for their detection as unusual stellar objects.
@@ 2 Numerical Framework
We aim to capture the details of the tidal disruption, which takes place over a few dynamical timescales of the star, as well as the remnant’s subsequent long-term evolution, lasting upwards of a few billion years. To address this challenge, we have combined simulations of TDEs using the adaptive mesh hydrodynamic code FLASH (B. Fryxell et al., 2000) paired with stellar evolution models from the Modules for Experiments in Stellar Astrophysics code (MESA version 10398, B. Paxton et al. 2011, 2013, 2015, 2018) that explore the subsequent long-term evolution of the disrupted star. The hydrodynamic simulations are explained in Section 2.1, while the stellar evolution setup is detailed in Section 2.2.
Table 1: Summary of TDE remnant model properties
Name Age $\beta$ $M_{i}$ $M_{f}$ Mass loss
(Gyr) ($M_{\odot}$) ($M_{\odot}$) (%)
$1M\beta 1$ 4.8 1 1 0.956 4.4
$1M\beta 1.5$ 4.8 1.5 1 0.708 29.2
$1M\beta 2$ 4.8 2 1 0.386 61.4
$3M\beta 4$ 0.3 4 3 1.06 64.7
@@ 2.1 Hydrodynamic FLASH models
In J. A. P. Law-Smith et al. (2020a), stellar models of varying initial masses and ages from MESA were mapped into 3D FLASH simulations (J. Guillochon et al., 2009; J. Guillochon & E. Ramirez-Ruiz, 2013), where they were tidally disrupted by a SMBH with mass $M_{\rm BH}=10^{6}\,M_{\odot}$. The results of these simulations are cataloged in the STARS Library (J. A. P. Law-Smith et al., 2020b). For our study of the long-term stellar evolution post-disruption, we select three disruptions at solar metallicity ($Z=0.0142$) of a star with initial mass $M_{i}=1\,M_{\odot}$ which were tidally disrupted at an age of 4.8 Gyr. Based on the known population of stars in the Galactic Center (R. D. Blum et al., 2003; R. Genzel et al., 2010; O. Pfuhl et al., 2011; J. R. Lu et al., 2013; R. Schödel et al., 2020), we choose to primarily work with solar-mass stars to strike a middle ground between massive stars, which are rare, and low-mass stars, which are more abundant but are often fully disrupted, leaving no remnant. We selected middle-aged main sequence stars (MAMS) as our progenitors, as that is the evolutionary phase where stars spend the majority of their lives. We also model a fourth remnant from an initially $M_{i}=3\,M_{\odot}$ progenitor, disrupted at an age of 0.3 Gyr (MAMS). The properties of each of the remnants modeled in this paper are summarized in Table 1.
@@
The models with a $M_{i}=1\,M_{\odot}$ progenitor underwent partial tidal disruption with varying values of $\beta=1$, 1.5, and 2. We define $\beta$ as the ratio of the tidal disruption radius ($R_{\tau}$) to the pericenter distance ($R_{p}$) – a dimensionless parameter that characterizes the strength of the disruption. For instance, when $\beta=1$ the pericenter distance is equal to the tidal radius, the point at which the SMBH’s tidal forces first begin to siphon mass from the star. Above $\beta=2$, the $M_{i}=1\,M_{\odot}$ progenitors from the STARS Library are usually fully, or nearly fully, destroyed (see Figure 5 in J. A. P. Law-Smith et al., 2020a). Thus for the $M_{i}=1\,M_{\odot}$ progenitor, we select $\beta$ values distributed between 1 and 2 to best capture the possible range of disruption remnants. This set of remnant models represents the range of outcomes for common stars that are likely to be disrupted, while also being likely to survive. The corresponding final masses of each TDE remnant model from $M_{i}=1\,M_{\odot}$ are $M_{f}=0.956,0.708,$ and $0.386\,~M_{\odot}$ for $\beta=1$, 1.5, and 2 respectively. We refer to these models by their $\beta$ value and progenitor mass, e.g., the $1M\beta 1$ remnant refers to the remnant disrupted at $\beta=1$ from a $M_{i}=1\,M_{\odot}$ progenitor.
@@
The $3M\beta 4$ remnant is partially tidally disrupted with $\beta=4$ from a progenitor with $M_{i}=3\,M_{\odot}$. Post-disruption, the remnant mass is $M_{f}=1.06\,M_{\odot}$, nearly the same as the initial mass of our other models. Its final mass is also very similar to the post-disruption mass of the $1M\beta 1$ remnant. As a result, the long-term evolution of the $3M\beta 4$ remnant provides an interesting contrast to that of similar-mass models with different initial conditions.
@@
Figure 1 shows snapshots of the four TDE remnants at the end of the FLASH simulation, which terminates roughly 80 $t_{\rm dyn}$ after pericenter, where $t_{\rm dyn}\equiv(GM_{\star}/R_{\star}^{3})^{-1/2}$ is the star’s dynamical time (J. A. P. Law-Smith et al., 2020a). The three panels on the left compare the three remnants from the same progenitor of $M_{i}=1\,M_{\odot}$. After the disruption, the least disrupted remnant ($1M\beta 1$) retains a dense, hot core with a high internal diffusion time, and a relatively short diffusion time in the envelope. This suggests the TDE primarily deposited energy close to the star’s surface without impacting much of the stellar interior. On the opposite end, the most disrupted remnant ($1M\beta 2$) is relatively cooler throughout with a diminished diffusion time at the core, indicating the TDE had a greater impact on the star’s internal structure. Energy deposited closer to the surface where the diffusion time is short will leak out first, powering the early-time luminosity of the remnant.
@@
On the right, we show snapshots from the $3M\beta 4$ remnant, whose radial extent is a factor of $\sim 2$ larger than the remnants from a $M_{i}=1\,M_{\odot}$ progenitor. As this progenitor was disrupted at a very large $\beta$, the remnant also exhibits a short diffusion time throughout the core that signifies deep energy deposition from the TDE. Given the increased radial scale of these snapshots, we can see that the $3M\beta 4$ remnant is much more diffuse, but still retains a relatively high core temperature due to its more massive progenitor.
@@
Figure 2: Luminosity, temperature and radius evolution of the three tidally disrupted remnants (red solid) with progenitor mass $M_{i}=1\,M_{\odot}$. These remnants are compared to mass-equivalent stars (blue dash-dotted) that evolve from the pre-main sequence (PMS) to the main sequence (MS) and up the red giant branch (RGB). The masses of the TDE remnants and their associated mass-equivalent stars are $0.956\,M_{\odot}$, $0.708\,M_{\odot}$, and $0.386\,M_{\odot}$ respectively from left to right. The start of core hydrogen fusion is denoted by star symbols for remnants and mass-equivalent stars. In addition, the evolution of an undisturbed $M=1\,M_{\odot}$ star (black dotted) is shown from the time of disruption onward in order to represent the comparative evolution of the progenitor had it not undergone the TDE. Immediately after disruption, the remnants appear larger and more luminous than the progenitor, then shrink over the next several million years. After reigniting core hydrogen fusion, they stabilize near the MS locations of their mass-equivalent stars.